# TAME: Temporal Audio-based Mamba for Enhanced Drone Trajectory Estimation and Classification

### v1

@@ 1 Introduction
Compact unmanned aerial vehicles (UAVs) are popular for their portability, ease of use, and affordability, making them valuable in fields such as transportation, photography, search and rescue. However, these same features also make them difficult to detect when used maliciously, posing threats to air traffic control and security and being exploited in warfare and border drug smuggling, as shown in Fig. 1.
Existing anti-UAV detection methods often rely on single-modality techniques [1, 2, 3, 4], such as radio signals, radar, or images, but these approaches can be error-prone, particularly due to limitations in radar cross section [5, 6, 7] or pixel size [8]. While some methods attempt to overcome these issues through multi-modal fusion [9, 10], combining inputs like multi-spectral images and audio data, they often rely on unrealistic assumptions [11], such as perfect targeting or high vantage points, making them less practical in real-world scenarios. Audio, however, presents a distinct advantage by providing reliable information about the direction [12, 13], type, and distance of sound sources [14, 15], which is largely unaffected by environmental conditions. This is particularly useful given the significant noise generated by UAVs, making them easier to detect through audio [9]. Recognizing these benefits, recent studies [10] have begun to advocate for integrating audio-based solutions to improve UAV detection accuracy.
In response to these challenges, we propose a novel anti-UAV model using audio, called Temporal Audio-based Mamba for Enhanced Drone Trajectory Estimation and Classification (TAME), which utilizes the selective state-space model (SSM) [16] to effectively process temporal sequences in audio data for UAV detection. This model enables omnidirectional 3D spatial area detection, addressing the limitations of current methods that are primarily focused on classification, tracking, and position estimation using visual data. While effective, traditional multi-modal fusion approaches demand high computational power, which limits their feasibility for mobile or wearable applications. By leveraging the efficiency of SSM, TAME offers a more practical solution for accurate and efficient UAV detection, especially in challenging real-world environments. Our contributions are summarized as follows:
@@ 2.3 Detection Head
| Modality | Network | Light | Dark | (%) | |||||||||
| APE | Acc(%) | APE | Acc(%) | ||||||||||
| Visual | VisualNet [14] | 0.24 | 0.39 | 0.32 | 0.65 | 99.7 | 1.98 | 6.10 | 8.13 | 11.45 | 11.3 | 6.05 | 55.5 |
| DarkNet [19] | 0.23 | 0.46 | 0.23 | 0.63 | 100 | 1.84 | 5.50 | 4.57 | 8.31 | 25.9 | 4.47 | 63.0 | |
| Audio-Visual | TalkNet [15] | 0.31 | 0.69 | 0.44 | 0.99 | 100 | 1.13 | 3.39 | 3.92 | 5.82 | 47.4 | 3.41 | 73.7 |
| AV-PED [14] | 0.31 | 0.50 | 0.59 | 0.97 | 98.5 | 0.58 | 1.54 | 2.26 | 3.13 | 80.7 | 2.01 | 89.6 | |
| AV-FDTI [10] | 0.13 | 0.31 | 0.38 | 0.58 | 99.6 | 0.35 | 1.06 | 1.10 | 1.89 | 88.3 | 1.24 | 94.0 | |
| Audio | AudioNet [14] | 0.60 | 1.76 | 1.59 | 2.80 | 79.8 | 0.60 | 1.76 | 1.59 | 2.80 | 79.8 | 2.80 | 79.8 |
| DroneChase [9] | 0.54 | 1.59 | 1.51 | 2.64 | 80.6 | 0.54 | 1.59 | 1.51 | 2.64 | 80.6 | 2.64 | 80.6 | |
| TAME(Ours) | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.55 | 98.0 | |

### v4

@@ 1 Introduction
Compact unmanned aerial vehicles (UAVs) [1] are popular for their portability, ease of use, and affordability, making them valuable in fields such as transportation [2, 3], photography [4, 5, 6, 7], search and rescue [8, 9, 10]. However, these same features also make them difficult to detect when used maliciously, posing threats to air traffic control and security and being exploited in warfare and border drug smuggling [11], as shown in Fig. 1.
Existing anti-UAV detection methods often rely on single-modality techniques [12, 13, 14, 15, 16, 17, 18], such as radio signals, radar, or images [19], but these approaches can be error-prone, particularly due to limitations in radar cross section [20, 21, 22] or pixel size [23]. While some methods attempt to overcome these issues through multi-modal fusion [24, 25, 26], combining inputs like multi-spectral images and audio data, they often rely on unrealistic assumptions [27], such as perfect targeting or high vantage points, making them less practical in real-world scenarios. Audio, however, presents a distinct advantage by providing reliable information about the direction [28, 29], classification [30, 31], and distance of sound sources [32, 33], which is largely unaffected by environmental conditions. This is particularly useful given the significant noise generated by UAVs, making them easier to detect through audio [24]. Recognizing these benefits, recent studies [25] have begun to advocate for integrating audio-based solutions to improve UAV detection accuracy.
In response to these challenges, we propose a novel anti-UAV model using audio, called Temporal Audio-based Mamba [30] for Enhanced Drone Trajectory Estimation and Classification (TAME), which utilizes the selective state-space model (SSM) [34] to effectively process temporal sequences in audio data for UAV detection. This model enables omnidirectional 3D spatial area detection, addressing the limitations of current methods that are primarily focused on classification, tracking, and position estimation using visual data. While effective, traditional multi-modal fusion approaches demand high computational power, which limits their feasibility for mobile or wearable applications. By leveraging the efficiency of SSM, TAME offers a more practical solution for accurate and efficient UAV detection, especially in challenging real-world environments. Our contributions are summarized as follows:
@@ 2.3 Detection Head
Modality | Network | Params | Light | Dark | (%) | |||||||||
(M) | APE | Acc(%) | APE | Acc(%) | ||||||||||
Visual | VisualNet [32] | 31.18 | 0.24 | 0.39 | 0.32 | 0.65 | 99.7 | 1.98 | 6.10 | 8.13 | 11.45 | 11.3 | 6.05 | 55.5 |
DarkNet [37] | 40.69 | 0.23 | 0.46 | 0.23 | 0.63 | 100 | 1.84 | 5.50 | 4.57 | 8.31 | 25.9 | 4.47 | 63.0 | |
Audio | TalkNet [33] | 15.80 | 0.31 | 0.69 | 0.44 | 0.99 | 100 | 1.13 | 3.39 | 3.92 | 5.82 | 47.4 | 3.41 | 73.7 |
Visual | AV-PED [32] | 32.36 | 0.31 | 0.50 | 0.59 | 0.97 | 98.5 | 0.58 | 1.54 | 2.26 | 3.13 | 80.7 | 2.01 | 89.6 |
Fusion | AV-FDTI [25] | 92.28 | 0.13 | 0.31 | 0.38 | 0.58 | 99.6 | 0.35 | 1.06 | 1.10 | 1.89 | 88.3 | 1.24 | 94.0 |
Audio | AudioNet [32] | 1.45 | 0.60 | 1.76 | 1.59 | 2.80 | 79.8 | 0.60 | 1.76 | 1.59 | 2.80 | 79.8 | 2.80 | 79.8 |
DroneChase [24] | 0.24 | 0.54 | 1.59 | 1.51 | 2.64 | 80.6 | 0.54 | 1.59 | 1.51 | 2.64 | 80.6 | 2.64 | 80.6 | |
TAME(Ours) | 6.61 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.11 | 0.30 | 0.34 | 0.55 | 98.0 | 0.55 | 98.0 |